Hi! I think this problems has to do with slope and linear equations. It would help if you could explain step by step what needs to be done and give the equation with the answer. Thanks so much! ** An antique clock is expected to be worth $350 after 2 years and $530 after 5 years. Write an equation to represent the expected value, and use it to determine the value the clock should be worth after 10 years. **
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

You have two points: (2,350) and (5,530)
You need an equation of the form y=mx+b
m=(530-350)/(5-2)
m=180/3
m=60
Using x=2, y=350, and m=(60) we solve for "b", as follows:
350=(60)(2)+b
b=230
Therefore the equation is y=(60)x+230
Finding the value of the clock after 10 years,
let x=10 to get y=60(10)+230
y=600+230
y=$830

Answer by checkley71(8403)

I DON'T HAVE THE ABILITY TO DRAW GRAPHS BUT YOUD PLOT YEARS ON THE X AXIS AND PRICE ON THE Y AXIS. YOUR FIRST POINT WOULD BE $350 AT YEAR 2 AND THE SECOND POINT AT $530 FOR YEAR 5. DRAW A STRAIGHT LINE THROUGH THESE POINTS AND CHECK THE PRICE AT THE 10 YEAR MARK WHICH WILL BE $830.

HERE'S THE ALGEBRAIC SOLUTION.
$530-$350=$180 THE INCREASE FOR THREE YEARS OR AN AVERAGE OF $60 PER YEAR FOR THE FIRST THREE YEARS. THEREFORE FOR THE NEXT 5 YEARS YOU WOULD EXPECT THE PRICE TO INCREASE 5X$60 OR $300 THEN THE PRICE WOULD BE $530+$300 OR $830
THE EQUATION IS ((530-350)/3)X(5)+530 OR (180/3)X(5)+530 OR (60X5)+530 OR
300+530=830